# Bulk–boundary correspondence from hyper-invariant tensor networks

### v1

@@ 3.1 Two point correlation function
Let us define a path between two boundary subsystems as a subset of tensors (together with its conjugates) from the tensor network, such that in each tensor two and only two of its indices, which are not adjacent, connect it with other tensors in the path or are the boundary indices corresponding to distinguished boundary subsystems. An example of a path is presented in Figure 5 (d). Note that one cannot reduce the path, by contracting it with its conjugate on all indices except the distinct one. It is so because each path tensor has at most two neighbouring indices which are neither distinct nor connecting it with other path tensors.
A special type of path is a geodesic – a path with all tensors connected in the same manner. One can think of geodesic as a path which in each node takes the same turn, as illustrated in Figure 4 (a). Notice that the definition of the path is slightly more general than geodesics. The example path which is not a geodesic is presented in 5 (d). It turns out that neither paths nor geodesics are as common as one would think.
@@
We prove this property by contradiction. Assume that for some boundary indices, there exist at least two paths $t_{1}$, $t_{2}$ connecting them. Let $n_{a}$ and $n_{b}$ be their last common nodes. Thus we may take two geodesics paths $l_{1}$, $l_{2}$ that start in $n_{a}$, follow the first step of $t_{1}$ and $t_{2}$ respectively and turn to each other. Since $t_{1}$ and $t_{2}$ join at $n_{b}$, the $l_{1}$ and $l_{2}$ must intersect in some node $n_{b}^{\prime}$ closer to $n_{a}$ than $n_{b}$. Therefore we constructed two geodesics on a discretized Poincare disc which intersects in two points, which is contrary to hyperbolic geometry. ∎
@@
Figure 5: Calculation of the 2-point correlation function, the tensors with distinct indices are denoted in red and the simplified tensors in grey. The entire network is presented in (a). In panel (b) the first two tensors, marked grey, are reduced, which ensures the reduction of almost all tensors from this layer (c). Similar procedure is then repeated until only the path between distinct indices remains (d) and the 2-point correlation function can be directly computed.
In the consecutive steps, one repeats these steps over the new boundary. This procedure stops only if all tensors are reduced, or none of the remaining tensors can be reduced. The latter implies that the remaining tensors have at least two non-neighbouring indices connected with other non-reduced tensors (or these indices are the distinct ones). Thus the boundaries of a set of unreduced tensors must be paths connecting distinct boundary indices. This means, due to the uniqueness of the path, that unreduced tensors are forming a path. The diagram of the above-discussed actions is presented in Figure 5.
@@
Note that not all subsystems are connected by a path within a network, thus not all local subsystems are correlated. It is, in a sense side effect of a discrete network construction, which imposes that not all subsystems can be connected by a discrete geodesic within such a network. On the other side, we stress once again that the defined notion of the path is more general than geodesic.
@@
Each simplified node on the path can be interpreted as a matrix acting on $8$ qudits. Such a matrix have an eigenvector $\sum_{\mathbb{i}}|\mathbb{i}\mathbb{i}\rangle$, with $\mathbb{i}=(i_{1},i_{2},i_{3},i_{4})$, to the eigenvalue $d^{5}=32$. Moreover, it is also an eigenvalue with the maximal module of this matrix.
@@
The first part of the Lemma can be established by a simple calculation, relying on properties of unitarity. To prove the second part we use the fact that the spectral radius of any matrix $r(A)$ is not larger than its operator norm, $r(A)\leq||A||=\max_{\mathbb{x}}||A|\mathbb{x}\rangle||/|||\,|\mathbb{x}\rangle||$ Conway2007 .
Let us consider any vector $|\mathbb{x}\rangle$ on which the reduced node acts. By using repetitively three properties
@@
one can show that the norm $||\,|\mathbb{x}\rangle||$ can increase at most $d^{5}=32$ times. ∎
@@
Assume the reduced path node has only one eigenvector to eigenvalue with maximal module $d^{5}=32$. Then by normalizing the nodes by this factor, in the limit of a large network, one obtains the desired decay behaviour of two-point correlation functions (2). Moreover, if the bulk operator has normalized trace and the path is a geodesic, the scaling dimension $\Delta$ is given by
@@
Consider one step of inflation – new layer addition, for the tensor network with a large number of layers. From one hand side, the number of tensors grows according to scaling factor $\mu=2+\sqrt{3}$ Central_charges_and_scaling , resulting in the same growth of distance between two points on the boundary. On the other hand new layer introduces two new nodes, each on opposite ends of the path, which lowers the correlation by a factor equal to the next to the leading eigenvalue squared. We neglect the terms with lower eigenvalues since they decay appropriately faster in a large number of layers limit. Thus we obtain the desired power-low decay of correlations (2). If the considered path of interest is a geodesic, it consists of exactly the same nodes, so we may compare correlations after the inflation step explicitly:
@@
In the presented model the discussed conformal field theory lies on a cylinder, with the constant time slice being a circle. Thus to be precise one should consider an exact correlation function for the theory mapped onto a cylinder Ginsparg:1988ui
@@
3.2 Three point correlation functions
@@
We note that the requirements for a non-vanishing three-point correlation function are, once again, a side effect of hyperbolic space discretization. In AdS space any two points can be connected via geodesic, thus the three-point correlation function would not disappear for changes in points’ configuration on the boundary.
@@
Through the proof, we use the fact that a tensor network can be placed on the tessellation of hyperbolic space, thus we may interpret the paths as polygonal chains. Moreover, we notice that each path lies between two geodesics tangent to any chosen connection in the path. Let us start by showing that there cannot exist $3$ paths connecting each pair of distinct indices. Lets consider three such paths $t_{1,2},t_{2,3},t_{1,3}$ intersecting in the nodes $n_{1},n_{2},n_{3}$.
One can replace paths $t_{2,3}$, $t_{1,3}$ by two geodesics $l_{1,3},l_{2,3}$, tangent to $t_{2,3}$, $t_{1,3}$, in line segments starting from points $n_{1}$ and $n_{2}$ respectively, and intersecting in the node $n_{3}^{\prime}$ on the same side as $n_{3}$ but not further from the path $l_{1,2}$. However, we can also replace the path $t_{1,2}$ by a geodesic $l_{1,2}$, which intersects with $l_{1,3}$ and $l_{2,3}$.
Since the hyperbolic plane was tiled with pentagons, the angle between line segments is $360^{\text{o}}/5=72^{\text{o}}$ and the angle between the geodesics cannot be smaller. Thus using geodesics $l_{1,2}$, $l_{2,3}$ and $l_{1,3}$ we constructed a triangle on the hyperbolic plane with a sum of inner angles larger than $180$ degrees, which is not possible.
@@
Figure 6: (a) Exemplary reduction of tensor network for three-point correlation function. Red tensors possess distinguished indices, while grey tensors can be reduced in the same manner as for 2-point correlation.
@@
With the same assumptions as in Corollary 1, the obtained three-point correlation function has the desired form (3) with the same scaling dimension as 2-point correlation functions     $\Delta=-\log_{\mu}\lambda_{2}$. The proportionality coefficient $C_{123}$ depends on the reduction of the bulk operator to the node in which geodesics intersect.
@@
Using the same method one can extend this analysis into a higher-order correlation function, however, the analysis becomes complicated due to all possible combinations in which paths between boundary indices may or may not intersect.
4 Numerical examples
In this Section, we present a numerical discussion of available scaling dimensions originating from two-point correlation functions between two boundary indices connected via geodesic.
@@ 4 Numerical examples
${U^{(2)}}_{ij,lk}=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&0&1&0\\ 0&\frac{1}{2}\left(1+e^{i\pi a}\right)&0&\frac{1}{2}\left(1-e^{i\pi a}\right)\\ 0&\frac{1}{2}\left(1-e^{i\pi a}\right)&0&\frac{1}{2}\left(1+e^{i\pi a}\right)\\ \end{array}\right)=CNOT^{a}\textit{SWAP},$ (13)
@@
The freedom to choose example $2$ ququart unitaries is much larger. However, for the sake of simplicity, we restricted ourselves to two cases. First was a product of two $CNOT^{b}$ gates, with power $b\in[0,1]$, interpolating between identity and $CNOT$ (see App. C), each acting on one frame’s qubit and one perfect tensor’s qubit, with the frame qubit being the control one. The second family of $2$ ququart unitaries was similar but this time we first on $CNOT$ onto a perfect tensor’s qubits and with $CNOT^{b}$ in the opposite direction, and power $b\in[0,1]$, effectively interpolating between $CNOT$ and $DCNOT$ (see App. C).
Thus, those two cases may be summarized as
@@
To probe the entire range of possible eigenvalues for reduced tensor nodes we considered random dual unitary and $2$ ququart unitary. Both unitaries were sampled with the Haar measure. However, to obtain the required properties of dual unitarity we used a Sinkhorn algorithm, presented example in 36_officers_of_Karol , using a random unitary sample as a starting point. The distribution of obtained eigenvalues and corresponding scaling dimensions are presented in Figure 8.
@@
As we mentioned in the introduction, the HaPPY tensor network does not reproduce the expected form of the two-point correlation functions, unless we interpret them as decaying infinitely fast (7). However, if we break some of the HaPPY’s symmetry, as we intentionally did in our construction, the scaling dimensions may obtain different nontrivial values. In our numerical study, we find the minimum and maximal values of the scaling dimension to be in the range $\Delta\in(0.858852,6.841722)$. Since the minimal value of the scaling dimension is greater than zero $\Delta>0$, that guarantees simulated conformal (scalar) field theory to be unitary Poland:2018epd . Moreover, since $\Delta=1$ is within the range on our model, we can simulate the primary fields $\phi_{\varepsilon}$ with conformal weights $(\frac{1}{2},\frac{1}{2})$, which corresponds, for instance, to the energy $\varepsilon$ in the Ising model discussed in Pfiefer2009 . This indicates that novel constructions based on the foundations of the HaPPY model could indeed be tuned to simulate wider ranges of properties for AdS/CFT correspondence.
@@
Similar results were earlier obtained by other approaches – see Steinberg2022 . However, up to our knowledge, none of the derivations concerns the networks simulating bulk-boundary correspondence, led to the expected form of the two-point correlation functions.
5 Concluding Remarks
@@
In this work, we constructed and investigated tensor networks inspired by the AdS/CFT correspondence. By combining perfect tensors HaPPy_code with hyper-invariant ones Evenbly2017 , as illustrated in Figure 3(b), we were able to benefit from the desired properties streaming from both of them. Using perfect tensors we reestablished a simple, yet elegant mapping from the bulk to the boundary Hilbert space, and thanks to the hyper-invariant frame we obtained nontrivial correlations at the boundary. Our construction is designed to resemble conformal field theory (CFT) on the boundary of a $\{5,4\}$ hyperbolic tessellation of the Poincaré disk.
@@
Inspired by the field-theoretic perspective, instead of discussing quantum states, we focused on mapping bulk operators to boundary ones by "sandwiching" them between the constructed tensor network and its conjugation. Such an approach allows us to achieve in a natural way the desired behaviour of two- and three-point correlation functions at the boundary, while considering the image of the bulk operators. This indicates that, together with hyperbolic geometry, we have established a bulk-boundary relation for HTN codes. The obtained result opens new avenues for studying HTN networks through explicit holographic-like mappings, which could be particularly interesting from the perspective of quantum error correction, as recently proposed for Evenbly codes Steinberg:2024ack .
@@
Due to flexibility in our construction, we managed to explore a wide range of scaling dimensions $\Delta$ within simple examples. This highlights the ability of the model to be tailored for desired behaviour, corresponding to certain conformal fields of interest. For instance, $\Delta=1$ corresponds to energy $\epsilon$ in the Ising model at the critical point Pfiefer2009 .
@@
The natural direction of further study would be to generalize the presented tensor network for different tessellations of Poincaré disc. Our methodology advocated here for perfect tensors with $6$ indices, can be applied to any other perfect tensor, with order greater or equal $6$ and "rotational symmetry". This approach may directly connect the curvature of AdS space, represented by the different tilings, with the properties of conformal field theory, expressed in correlation functions. The question, whether the proposed approach enables one to obtain a discrete counterpart of the Virasoro algebra remains open. Such a task, already achieved for quantum CFT Jahn2022 , allows one to recover the energy-momentum tensor of boundary conformal field theory. Addressing this problem can provide insights into the gravitational aspects of a compatible bulk theory, thereby deepening the connection between holography and tensor networks. This direction of research could substantially advance comprehension of the interplay between quantum gravity and quantum information theory.
Acknowledgements.
@@
It is a pleasure to thank Mario Flory, Felix Huber, Romuald Janik, Gerard Munne, Zahra Raissi and Marcin Rudziński for their fruitful remarks and suggestions. This work was supported by the National Science Centre, Poland, under the contract number 2021/03/Y/ST2/00193 within the DQuant QuantERA II project that has received funding from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No 101017733.
Appendix A Building blocks of the construction proposed
@@
In this Appendix, we present a complete and explicit formula for the construction of the proposed hyper-invariant tensor, following the steps from Section 2.
@@
The first ingredient of our recipe is a perfect tensor $T$ of rank $6$ with cyclic symmetry on all except first index $T_{i,jklnm}=T_{i,njklm}$. Although we didn’t manage to obtain such a form for the smallest possible local dimension $2$, the construction of a clear example can be provided already for the perfect tensor of the local dimension $D=4$. As perfect tensor we choose a construction of three orthogonal Latin cubes Goyeneche:2015fda presented in Figure 9. Latin cube is an example of combinatorial design, which can be represented as a cube $L_{a,b,c}$ with discrete values such that each hyper-row of the cube is filled with numbers $1,\cdots,D$ without repetition. Orthogonality of Latin cubes can be expressed as an orthogonality between their consecutive layers treated as Latin squares, see ex. bistron2023genuinely .
@@
Figure 9: Three orthogonal Latin cubes of dimension $D=4$ corresponding to an AME state of six ququarts and the perfect tensor with six indices running from one to four (16). Figure borrowed from Goyeneche:2015fda
@@
The next step of our recipe is the construction of a hyper-invariant frame $F$. Similarity as Evenbly2017 , we build then with bipartite double unitary matrices, which satisfy not one, but two orthogonality relations
@@
Appendix B On hyper-invariant tensor networks
For the convenience of the readers in this Appendix we present the definition of hyperinvariant tensor network and the Evenbly code, consistent with the notation from the rest of the work, following Evenbly2017 ; Steinberg:2024ack .
@@ Appendix B On hyper-invariant tensor networks
An Evenbly code, called also hyper-invariant code Steinberg2023 , is a modification of the hyper-invariant tensor network, in which tensors $A$ have an additional index $i_{0}$ used to encode information from the bulk. The following definition is a slightly relaxed version of the one form Steinberg:2024ack .
@@
Appendix C On bipartite unitary matrices
Dual unitary matrices BKP19 ; BRRL24 are a building stone of the proposed construction (18). In this Appendix, we discuss the bipartite unitary matrices of order $d^{2}$, demonstrating particular features of the subset of dual unitary matrices. Most of the discussion concerns two-qubit gates, $d^{2}=4$, however, some of the below-mentioned properties hold also for arbitrary local dimension $d$.
Any two-qubit unitary matrix $U_{AB}\in U(4)$ can be conveniently represented in its Cartan form 2_qubit_gate ; HVC01 ,
@@ Appendix C On bipartite unitary matrices
Here $\sigma_{1},\sigma_{2}$ and $\sigma_{3}$, represent Pauli matrices, while the phases $\alpha_{i}\in[0,2\pi]$ characterizing interaction strength. The choice of one-qubit gates does not affect the orthogonality relations (17), thus we may focus directly on the interaction part. Moreover, to avoid over-parametrization, one can restrict the values of parameters $\pi/4\geq\alpha_{1}\geq\alpha_{2}\geq\alpha_{3}\geq 0$, effectively creating a simplex, which forms half of the Weyl chamber ZVSW03 ; Weyl_chamber .
@@
$Id=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\\ \end{array}\right),CNOT=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\\ \end{array}\right),DCNOT=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&0&0&1\\ 0&1&0&0\\ 0&0&1&0\\ \end{array}\right),\textit{SWAP}=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&0&1&0\\ 0&1&0&0\\ 0&0&0&1\\ \end{array}\right),$

### v2

@@ III.1 Two point correlation function
Let us define a path between two boundary subsystems as a subset of tensors (together with its conjugates) from the tensor network, such that in each tensor two and only two of its indices, which are not adjacent, connect it with other tensors in the path or are the boundary indices corresponding to distinguished boundary subsystems. An example of a path is presented in Figure 6 (d). Note that one cannot reduce the path, by contracting it with its conjugate on all indices except the distinct one. It is so because each path tensor has at most two neighboring indices which are neither distinct nor connecting it with other path tensors.
A special type of path is a geodesic path, where for each path’s node one edge of the corresponding tile belongs to the same geodesic in AdS space. One can think of a geodesic path as a path which in each node takes the same turn, as illustrated in Figure 5 (a). Notice that the definition of the path is slightly more general than the geodesic path. The example path which is not a geodesic path is presented in 6 (d). It turns out that paths are as common as one would think.
@@
We prove this property by contradiction. Assume that for some boundary indices, there exist at least two paths $t_{1}$, $t_{2}$ connecting them. Let $n_{a}$ and $n_{b}$ be their last common nodes. Thus we may take two geodesics $l_{1}$, $l_{2}$ that include edges of $n_{a}$, and of the next node along $t_{1}$ and $t_{2}$ respectively, that goes between $t_{1}$ and $t_{2}$. Since $t_{1}$ and $t_{2}$ join at $n_{b}$, and each path could not cross a geodesic (it could be at most "tangent" to it), the $l_{1}$ and $l_{2}$ must intersect at the vertex of some node $n_{b}^{\prime}$ closer to $n_{a}$ than $n_{b}$. Therefore we constructed two geodesics on a Poincare disc which intersects in two points, which is contrary to hyperbolic geometry. ∎
@@
Figure 6: (a) Calculation of the 2-point correlation function between the points at the boundary with corresponding tensors marked red. (b) two tensors, marked in gray, are reduced, which ensures reduction of other tensors from this layer (c). Similar procedure is repeated until only the path between distinct indices (red symbols) remains. (d) 2-point correlation function can be directly computed.
In the consecutive steps, one repeats these steps over the new boundary. This procedure stops only if all tensors are reduced, or none of the remaining tensors can be reduced. The latter implies that the remaining tensors have at least two non-neighboring indices connected with other non-reduced tensors (or these indices are the distinct ones). Thus the boundaries of a set of unreduced tensors must be paths connecting distinct boundary indices. This means, due to the uniqueness of the path, that unreduced tensors are forming a path. The diagram of the above-discussed actions is presented in Figure 6.
@@
Note that not all subsystems are connected by a path within a network, thus not all local subsystems are correlated. It is, in a sense side effect of a discrete network construction, which imposes that not all subsystems can be connected by a geodesic path within such a network. On the other side, we stress once again that the defined notion of the path is more general than the geodesic path.
@@
Each simplified node on the path can be interpreted as a matrix acting on $8$ qudits. Such a matrix have an eigenvector $\sum_{\mathbf{i}}|\mathbf{i}\mathbf{i}\rangle$, with $\mathbf{i}=(i_{1},i_{2},i_{3},i_{4})$, to the eigenvalue $d^{5}=32$. Moreover, it is also an eigenvalue with the maximal module of this matrix.
@@
The first part of the Lemma can be established by a simple calculation, relying on properties of unitarity. To prove the second part we use the fact that the spectral radius of any matrix $r(A)$ is not larger than its operator norm, $r(A)\leq||A||=\max_{\mathbf{x}}||A|\mathbf{x}\rangle||/|||\,|\mathbf{x}\rangle||$ [33].
Let us consider any vector $|\mathbf{x}\rangle$ on which the reduced node acts. By using repetitively three properties
@@
one can show that the norm $||\,|\mathbf{x}\rangle||$ can increase at most $d^{5}=32$ times. ∎
@@
Assume the reduced path node has only one eigenvector to eigenvalue with maximal module $d^{5}=32$. Then by normalizing the nodes by this factor, in the limit of a large network, one obtains the desired decay behavior of two-point correlation functions (2). Moreover, if the bulk operator has normalized trace and the path is a geodesic, the scaling dimension $\Delta$ is given by
@@
Consider one step of inflation – new layer addition, for the tensor network with a large number of layers. From one hand side, the number of tensors grows according to scaling factor $\mu=2+\sqrt{3}$ [31], resulting in the same growth of distance between two points on the boundary. On the other hand new layer introduces two new nodes, each on opposite ends of the path, which lowers the correlation by a factor equal to the next to the leading eigenvalue squared. We neglect the terms with lower eigenvalues since they decay appropriately faster in a large number of layers limit. Thus we obtain the desired power-low decay of correlations (2). If the considered path of interest is a geodesic path, it consists of exactly the same nodes, so we may compare correlations after the inflation step explicitly:
@@
In the presented model the discussed conformal field theory lies on a cylinder, with the constant time slice being a circle. Thus to be precise one should consider an exact correlation function for the theory mapped onto a cylinder [32]
@@
III.2 Three point correlation functions
@@
We note that the requirements for a non-vanishing three-point correlation function are, once again, a side effect of hyperbolic space discretization. In AdS space any two points can be connected via geodesic, thus the three-point correlation function would not disappear for any points’ configuration on the boundary.
@@
Let us start by showing that there cannot exist $3$ paths connecting each pair of distinct indices. Lets consider three such paths $t_{1,2},t_{2,3},t_{1,3}$, and mark their last common nodes $n_{1},n_{2},n_{3}$.
Similarly as in the proof of the Lemma 1 one can replace paths $t_{1,3}$, $t_{2,3}$ by two geodesics $l_{1,3},l_{2,3}$, that contains the edges of $n_{1}$ and $n_{2}$, "follows" $t_{2,3}$ and $t_{1,3}$, and are between those paths. Because paths $t_{1,3}$ and $t_{2,3}$ cannot cross geodesics $l_{1,3}$ and $l_{2,3}$, those geodesic must intersecting in the vertex of some node $n_{3}^{\prime}$ on the same side as $n_{3}$ but not further from the path $l_{1,2}$. However, in a similar manner, we can also replace the path $t_{1,2}$ with a geodesic $l_{1,2}$, that intersects with $l_{1,3}$ and $l_{2,3}$.
In the used tiling of the AdS space four pentagons meet in each vertex, so the angles between intersecting geodesic are $360^{\text{o}}/4=90^{\text{o}}$. Thus using geodesics $l_{1,2}$, $l_{2,3}$ and $l_{1,3}$ we constructed a triangle on the hyperbolic plane with a sum of inner angles larger than $180$ degrees, which gives a contradiction.
@@
Figure 7: (a) Exemplary reduction of tensor network for three-point correlation function. Red tensors possess distinguished indices, while gray tensors can be reduced in the same manner as for 2-point correlation.
@@
With the same assumptions as in Corollary 1, the obtained three-point correlation function has the desired form (3) with the same scaling dimension as 2-point correlation functions     $\Delta=-\log_{\mu}\lambda_{2}$. The proportionality coefficient $C_{123}$ depends on the reduction of the bulk operator to the node in which paths intersect.
@@
III.3 Higher correlation functions
Using the same methods one can extend analysis into other correlation functions. However, the higher the order of the correlation function, the more complex analysis becomes. New phenomenon that appears at the level of the four-point correlation function, is a set of nodes that can not be reduced, even though they do not belong to the paths connecting distinct boundary indices. Such an effect does not occur for three-point correlation functions, since it would lead to the construction of a triangle with three inner angles equal to $\pi/2$ on the hyperbolic plane – consult proof of Theorem 2.
However, while studying the four-point correlation function one can consider a "rectangle" of nodes of shape $2\times k$, as a segment of two parallel neighboring paths, tangent to a geodesic between them. Such a set of nodes cannot be reduced, for example, if two shorter sides of this "rectangle" are contained in paths that connect distinct boundary indices, see Fig 8.
Figure 8: Exemplary reduction of tensor network for correlation function between four boundary indices with corresponding tensors marked red: scenario in which some non-reduced tensors do not belong to paths between distinct boundary indices.
For higher point correlation functions such cliques of non-reduced nodes become larger and more complicated. Similar phenomena were observed in the reductions of hyper-invariant codes [23].
Despite those complications, the argument regarding the decay of correlation functions still holds. Adding one more layer to the tensor network will extend each path to a distinct boundary index by one node, resulting, in a large network limit, in the appropriate scaling, analogously as in Corollary 2. Furthermore, convoluted formulas for higher point correlation functions, which cannot be reduced into lower order correlations, display a highly interactive nature of speculated field theory simulated at the boundary.
IV Numerical examples
In this Section, we present a numerical discussion of available scaling dimensions originating from two-point correlation functions between two boundary indices connected via a geodesic path.
@@ IV Numerical examples
${U^{(2)}}_{ij,lk}=\frac{1}{2}\left(\begin{array}[]{cccc}1&0&0&0\\ 0&0&2&0\\ 0&1+e^{i\pi a}&0&1-e^{i\pi a}\\ 0&1-e^{i\pi a}&0&1+e^{i\pi a}\\ \end{array}\right)=CNOT^{a}\textit{SWAP},$ (13)
@@
The freedom to choose example $2$ ququart unitaries is much larger. However, for the sake of simplicity, we restricted ourselves to two cases. First was a product of two $CNOT^{b}$ gates, with power $b\in[0,1]$, interpolating between identity and $CNOT$ (see App. C), each acting on one frame’s qubit and one perfect tensor’s qubit, with the frame qubit being the control one. The second family of $2$ ququart unitaries was similar but this time we first on $CNOT$ onto a perfect tensor’s qubits and with $CNOT^{b}$ in the opposite direction, and power $b\in[0,1]$, effectively interpolating between $CNOT$ and $DCNOT$ (see App.C). Thus, those two cases may be summarized as
@@
To probe the entire range of possible eigenvalues for reduced tensor nodes we considered random dual unitary and $2$ ququart unitary. Both unitaries were sampled with the Haar measure. However, to obtain the required properties of dual unitarity we used a Sinkhorn algorithm, presented example in [34], using a random unitary sample as a starting point. The distribution of obtained eigenvalues and corresponding scaling dimensions are presented in Figure 10.
@@
As we mentioned in the introduction, the HaPPY tensor network does not reproduce the expected form of the two-point correlation functions, unless we interpret them as decaying infinitely fast (7). However, if we break some of the HaPPY’s symmetry, as we intentionally did in our construction, the scaling dimensions may obtain different nontrivial values. In our numerical study, we find the minimum and maximal values of the scaling dimension to be in the range $\Delta\in(0.858852,6.841722)$. Since the minimal value of the scaling dimension is greater than zero $\Delta>0$, that guarantees simulated conformal (scalar) field theory to be unitary [35]. Moreover, since $\Delta=1$ is within the range on our model, we can simulate the primary fields $\phi_{\varepsilon}$ with conformal weights $(\frac{1}{2},\frac{1}{2})$, which corresponds, for instance, to the energy $\varepsilon$ in the Ising model discussed in [13]. This indicates that novel constructions based on the foundations of the HaPPY model could indeed be tuned to simulate wider ranges of properties for AdS/CFT correspondence.
@@
Similar results were earlier obtained by other approaches – see [17]. However, up to our knowledge, none of the derivations concerns the networks simulating bulk-boundary correspondence, led to the expected form of the two-point correlation functions.
V Concluding Remarks
@@
In this work, we constructed and investigated tensor networks inspired by the AdS/CFT correspondence. By combining perfect tensors [4] with hyper-invariant ones [15], as illustrated in Figure 3(b), we were able to benefit from the desired properties streaming from both of them. Using perfect tensors we reestablished a simple, yet elegant mapping from the bulk to the boundary Hilbert space, and thanks to the hyper-invariant frame we obtained nontrivial correlations at the boundary. Our construction is designed to resemble conformal field theory (CFT) on the boundary of a $\{5,4\}$ hyperbolic tessellation of the Poincaré disk.
@@
Inspired by the field-theoretic perspective, instead of discussing quantum states, we focused on mapping bulk operators to boundary ones by "sandwiching" them between the constructed tensor network and its conjugation. Such an approach allows us to achieve the desired behavior of two- and three-point correlation functions at the boundary, while considering the image of the bulk operators. This indicates that, together with hyperbolic geometry, we have established a bulk-boundary relation for the presented tensor network. The obtained model presents a new perspective for studying hyper-invariant tensor networks through explicit holographic-like mappings, which could be interesting from the viewpoint of quantum error correction, as recently proposed for Evenbly codes [26].
@@
Due to flexibility in our construction, we managed to explore a wide range of scaling dimensions $\Delta$ within simple examples. This highlights the ability of the model to be tailored for desired behavior, corresponding to certain conformal fields of interest. For instance, $\Delta=1$ corresponds to energy $\epsilon$ in the Ising model at the critical point [13].
@@
The natural direction of further study would be to generalize the presented tensor network for different tessellations of Poincaré disc. Our methodology advocated here for perfect tensors with $6$ indices, can be applied to any other perfect tensor, with order greater or equal $6$ and "rotational symmetry". This approach may directly connect the curvature of AdS space, represented by the different tilings, with the properties of conformal field theory, expressed in correlation functions. The question, whether the proposed approach enables one to obtain a discrete counterpart of the Virasoro algebra remains open. Such a task, already achieved for quantum CFT [16], allows one to recover the energy-momentum tensor of boundary conformal field theory. Addressing this problem can provide insights into the gravitational aspects of a compatible bulk theory, thereby deepening the connection between holography and tensor networks. This direction of research could substantially advance comprehension of the interplay between quantum gravity and quantum information theory.
Acknowlegments
@@
It is a pleasure to thank Mario Flory, Felix Huber, Romuald Janik, Gerard Munné, Zahra Raissi and Marcin Rudziński for their fruitful remarks and suggestions. This work was supported by the National Science Centre, Poland, under the contract number 2021/03/Y/ST2/00193 within the DQuant QuantERA II project that has received funding from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No 101017733.
Appendix A Appendix A. Building blocks of the construction proposed
@@
In this Appendix, we present a complete and explicit formula for the construction of the proposed hyper-invariant tensor, following the steps presented in the main part of the work.
@@
The first ingredient of our recipe is a perfect tensor $T$ of rank $6$ with cyclic symmetry on all except first index $T_{i,jklnm}=T_{i,njklm}$. Although we didn’t manage to obtain such a form for the smallest possible local dimension $2$, the construction of a clear example can be provided already for the perfect tensor of the local dimension $D=4$. As perfect tensor we choose a construction of three orthogonal Latin cubes [6] presented in Figure 11. Latin cube is an example of a combinatorial design, which can be represented as a cube $L_{a,b,c}$ with discrete values such that each hyper-row of the cube is filled with numbers $1,\cdots,D$ without repetition. Orthogonality of Latin cubes can be expressed as an orthogonality between their consecutive layers treated as Latin squares [36].
@@
Figure 11: Three orthogonal Latin cubes of dimension $D=4$ corresponding to an AME state of six ququarts and the perfect tensor with six indices running from one to four (16). Figure borrowed from [6]
@@
The next step of our recipe is the construction of a hyper-invariant frame $F$. Similarity as [15], we build then with bipartite double unitary matrices, which satisfy not one, but two orthogonality relations
@@
Appendix B Appendix B. On hyper-invariant tensor networks
For the convenience of the readers in this Appendix we present the definition of hyperinvariant tensor network and the Evenbly code, consistent with the notation from the rest of the work, following [15, 26].
@@
An Evenbly code, called also hyper-invariant code [23], is a modification of the hyper-invariant tensor network, in which tensors $A$ have an additional index $i_{0}$ used to encode information from the bulk. The following definition is a slightly relaxed version of the one form [26].
@@
Appendix C Appendix C. On bipartite unitary matrices
Dual unitary matrices [19, 20] are a building stone of the proposed construction (18). In this Appendix, we discuss the bipartite unitary matrices of order $d^{2}$, demonstrating particular features of the subset of dual unitary matrices. Most of the discussion concerns two-qubit gates, $d^{2}=4$, however, some of the below-mentioned properties hold also for arbitrary local dimension $d$.
Any two-qubit unitary matrix $U_{AB}\in U(4)$ can be conveniently represented in its Cartan form [37, 38],
@@
Here $\sigma_{1},\sigma_{2}$ and $\sigma_{3}$, represent Pauli matrices, while the phases $\alpha_{i}\in[0,2\pi]$ characterizing interaction strength. The choice of one-qubit gates does not affect the orthogonality relations (17), thus we may focus directly on the interaction part. Moreover, to avoid over-parametrization, one can restrict the values of parameters $\pi/4\geq\alpha_{1}\geq\alpha_{2}\geq\alpha_{3}\geq 0$, effectively creating a simplex, which forms half of the Weyl chamber [39, 40].
@@
$Id=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\\ \end{array}\right),CNOT=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\\ \end{array}\right),DCNOT=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&0&0&1\\ 0&1&0&0\\ 0&0&1&0\\ \end{array}\right),\textit{SWAP}=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&0&1&0\\ 0&1&0&0\\ 0&0&0&1\\ \end{array}\right),$ (20)